Hi May I have some help with this equation.
Solve the equation
x + 4| = |3x + 6|

Answer by stanbon(75887)

|x + 4| = |3x + 6|
-----------------------
x+4 = 3x+6
2x=-2
x=-1
----------
x+4 = -(3x+6)
x+4 = -3x-6
-4x = 10
x = -5/2
--------------
Cheers,
Stan H.